Find the following product
(2y+2)^2

Found 2 solutions by checkley77, jaya:
Answer by checkley77(12844)

(2y+2)^2
2y+2
2y+2 now multiply
----------------
4y^2+4y+4y+4
4y^2+8y+4
4(y^2+2y+1) ans.

Answer by jaya(3)

(2y+2)^2: To solve this problem we can use the formula:
(a+b)^2=a^2+2.a.b+b^2
=>(2y+2)^2=(2y)^2+2.2y.2+2^2
=> = 4y^2+8y+4